Operations Report – February 1st

Crew 220 Operations Report 01FEB2020

Sol: 13

Name of person filing report: Matt Wise

Non-nominal systems: Primary Generator, Hab lower deck 1 fluorescent light ballast (MDRS team in process of upgrade). Hot water heater has been repaired

Generator: Primary is inoperable. Rental was run-the following information is for the rental. MDRS team worked on primary generator yesterday making several startup attempts

Hours run: 12.5
From what time last night: 19:00

To what time this morning: 0730 (still off, SOC 91%, nice sunny day)

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 80%

Diesel Reading – 1/4

Station Propane Reading- 72%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50

Water Meter (units)- 0148719,2

Water (static tank) (gal) – 6 (MDRS team refilling in progress)

Static to loft Pump used – Yes 75 gal

Water in GreenHab (gal): 92.14

Water in Science Dome (gal): 0

Toilet Tank Emptied – No

Deimos rover used: Still in the workshop


Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: No


Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used – No

Starting Hours: 128.7

Beginning Charge: 100%

Ending Hours: 128.7

Ending Charge: 100%

Opportunity Rover used: Still in workshop


Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: No

Starting Hours: 134.7

Beginning Charge: 100%

Ending Hours: 134.7

Ending Charge: 100%

Notes on rovers: Opportunity and Deimos off-site for maintenance. Should be operational within a few sols

ATVs Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? Checked and added as necessary

# hours the ATV’s were used today: 0

Notes on ATVs: Maintenance performed on all. Honda Non-operational, 350-1 operational, 350-2 Non-operational, 350-3 Fully operational

Habcar used and why, where? yes; water. Major cleaning/maintenance by MDRS team
Trans fluid: ok
Coolant: 2/3 OK
Break Fluid: Full OK
Oil: OK
Power Steering Fluid: Full OK
Washer Fluid: Replacement part to be ordered

CrewCar used and why, where? Yes. Supplies brought to MDRS station. Incoming crew transported to station
Trans Fluid: OK
Coolant: 3/4 OK
Break Fluid: Full OK
Oil: OK
Power Steering Fluid: Full OK
Washer Fluid: Min. OK

General Notes and Comments: Cleaning and maintenance performed

Summary of the internet: Nominal. New TP Link wifi router donated and configured by Crew 220 Engineer. Username and password are both ‘admin’ for config interface. Original Netgear router is inside the tv cabinet in the grey carton.

Summary of Suit and Radios: 5 radios malfunctioning
Suits Nominal

Summary of Hab: Main fluorescent lights in Hab lower deck are being upgraded by MDRS team. New stove installed. New cameras installed and connected to existing station TV. Water heater repaired and functioning.
New Vizio 52″ flat panel TV donated to MDRS by crew 220 Engineer (wall mount included if you need it).

Summary of Science Dome operations: Nominal.

Summary of Ram operations: Nominal. MDRS team began work to upgrade exterior outlets

Summary of any observatory issues: Nominal

Summary of Health and Safety issues: MDRS crew all healthy

This is the last ops report for Crew 220. Thanks for everything!

Copyright © The Mars Society. All rights reserved. | Main Site